The initial kinetic energy of a block moving on a horizontal floor is 48 J. A constant frictional force acts on the block bringing it to rest over a distance of 2 m. What is the frictional force on the block?
Answer #1

work done = change in kinetic energy

W = Final kinetic energy - initial kinetic energy

= 0 - 48 = -48J

work done by frictional force = F*d*cos a

-48 = F*2*cos 180

F= 24 N

so the frictional force on the block is 24 N